OpenStudy (anonymous):

A plane intersects a sphere to form a circle. The distance from the center of the circle to the center of the sphere is 12 inches. If the area of the circle formed is 1225(pi)in.^2, what is the volume of the sphere?

mathslover (mathslover):

@ITTCandidate welcome to openstudy. For solving this question, you will first require to know the formula for finding the Volume of the sphere. Do you know what is that formula?

OpenStudy (anonymous):

I do, but I do not have the information I need to get that far.

OpenStudy (anonymous):

all the info I have is for the circle where the plane passes through

OpenStudy (anonymous):

there is a screen shot of the problem.

OpenStudy (anonymous):

I have been wracking my brain at this problem for 2 hours now. very discouraged.

OpenStudy (jack1):

use pythagoras rule to get the radius of the sphere, you have two sides of a right angled triangle there

OpenStudy (anonymous):

I know that the radius of the circle is 19.75(pi)

OpenStudy (anonymous):

Jack, I only have 1 leg, not 2, and not for the sphere

OpenStudy (jack1):

from the picture, that you have is a distance of 12 inches of the center of the circle to the center of the sphere, you have the radius of the circle, and from anywhere along the circumference of the circle in a direct line back to the center of the sphere is the radius of the sphere h^2 = a^2 + b^2 so calculate for h

OpenStudy (jack1):

I worked out that the radius of the circle is somewhere around the 35.43 mark, is that what you got?

OpenStudy (anonymous):

alright, so what I come up with is: 12^2+19.75(pi)^2=18.41 but the sqrt of that is only 4.29 which just can't be right... right?

OpenStudy (anonymous):

the radius I got was 19.75(pi)

OpenStudy (jack1):

ok so let's go back to your circles' radius 19 x 3.14 = just under 60... that seems way to big

OpenStudy (anonymous):

my answer is supposed to include the term "pi"

OpenStudy (jack1):

Area = pi x r^2 = 1255 pi so r = sqrt 1255 = 35.42598 ... you can put your final answer of volume in terms of pi, until then use normal numbers otherwise you'll confuse yourself

OpenStudy (anonymous):

the formula I used (perhaps incorrectly) was A=pi*r^2 which given my information was: 1225(pi)=pi*r^2

OpenStudy (jack1):

do you follow what I did though...? 1255 pi = pi x r^2 so 1255 = r^2 so r = sqrt 1255

OpenStudy (anonymous):

1225*3.14=3846.5

OpenStudy (jack1):

ok so Area = 1255 x pi = 3846.5 and Area = pi x r^2 so 3846.5 = pi x r^2 so 3846.6 / pi = r^2 so 1255 = r^2 35.43 = r

OpenStudy (anonymous):

well, crap... I've been looking at it wrong for the past couple hours... <frustrated>

OpenStudy (phi):

\[ \pi r^2 = 1225 \pi\] \[ r= \sqrt{1225}= 35 \]

OpenStudy (jack1):

ahhh, my bad, i thought it was 1255

OpenStudy (jack1):

so there we go, we've got our values of 12 for one side of the triangle, and 35 for the other, you find the value for h now and from there the volume post it back here if you don't mind, would love to see if you get the right answer now...?

OpenStudy (anonymous):

so the radius of the sphere is 37.41? does that sound right?

OpenStudy (jack1):

sorry, as @phi said the radius of the circle is 35, as the area was 1225 pi not 1255 that i was writing, my bad

OpenStudy (jack1):

so from that, you should get: \[h ^{2} = 12^{2}+35^{2}\]

OpenStudy (jack1):

so h^2 = 1369, and from that the sq root is 37 even.

OpenStudy (jack1):

so you all good with the above...? and what did you get for volume?

OpenStudy (anonymous):

not sure how to put it back into terms of (pi) but I got 125000.71in^2

OpenStudy (jack1):

... what equation are you using for volume?

OpenStudy (anonymous):

(4/3)pi*r^3

OpenStudy (anonymous):

and each time I enter it I'm getting a different answer

OpenStudy (anonymous):

now I get 212067.2266667

OpenStudy (jack1):

ok so 37^3 = 50653... yeah?

OpenStudy (anonymous):

ya

OpenStudy (jack1):

50653 x 4/3 = 67573.3... we matching so far?

OpenStudy (anonymous):

yes

OpenStudy (jack1):

so your final answer should be 212174.8... but we're leaving it in terms of pi so 67573.3 pi = volume

OpenStudy (anonymous):

no... 212067.226666

OpenStudy (anonymous):

I"m truncating pi to 3.14

OpenStudy (jack1):

ok, but i thought you wanted your answer in terms of pi...?

OpenStudy (anonymous):

correct

OpenStudy (jack1):

otherwise yeah, 212067

OpenStudy (jack1):

either way, think we're good now, slaters dude

OpenStudy (anonymous):

thank you!!!